# Deep learning-driven likelihood-free parameter inference for 21-cm forest observations

### v1

@@ Abstract
The hyperfine structure absorption lines of neutral hydrogen in spectra of high-redshift radio sources, known collectively as the 21-cm forest, have been demonstrated as a sensitive probe to the small-scale structures governed by the dark matter (DM) properties, as well as the thermal history of the intergalactic medium regulated by the first galaxies during the epoch of reionization. By statistically analyzing these spectral features, the one-dimensional (1D) power spectrum of the 21-cm forest can effectively break the parameter degeneracies and constrain the properties of both DM and the first galaxies. However, conventional parameter inference methods face challenges due to computationally expensive simulations for 21-cm forest and the non-Gaussian signal characteristics. To address these issues, we introduce generative normalizing flows for data augmentation and inference normalizing flows for parameters estimation. This approach efficiently estimates parameters from minimally simulated datasets with non-Gaussian signals. Using simulated data from the upcoming Square Kilometre Array (SKA), we demonstrate the ability of the deep learning-driven likelihood-free approach to generate accurate posterior distributions, providing a robust and efficient tool for probing DM and the cosmic heating history using the 1D power spectrum of 21-cm forest in the era of SKA. This methodology is adaptable for scientific analyses with other unevenly distributed data.
@@
Key Laboratory of Cosmology and Astrophysics (Liaoning) & College of Sciences, Northeastern University, Shenyang 110819, China
@@
Key Laboratory of Data Analytics and Optimization for Smart Industry (Ministry of Education), Northeastern University, Shenyang 110819, China
The hyperfine structure transition of neutral hydrogen atoms from the epoch of reionization (EoR) creates narrow absorption lines on spectra of high-redshift radio bright sources (e.g. radio-loud quasars), with a rest-frame wavelength of 21 cm. The 21-cm absorption lines from various structures at different distances along the line of sight are redshifted to different frequencies, giving rise to forest-like features on the spectra, which is called the 21-cm forest 1, 2, 3, 4, 5, 6 in analogy to the Lyman-$\alpha$ (Ly-$\alpha$) forest at lower redshifts. The 21-cm forest has been demonstrated to be a sensitive probe of the cosmic heating history 4, 5. After the formation of the first galaxies, the generated light not only ionizes hydrogen atoms but also heats the gas simultaneously (with X-rays being the primary heating source). As the spin temperature of the gas increases, the absorption at 21 cm weakens (reducing the optical depth), thus causing the 21-cm forest signal to become weaker. In principle, the strength of the signal can be used to measure the cosmic heating history, thereby gaining insight into the properties of the first galaxies.
On the other hand, understanding the properties of dark matter (DM) hinges crucially on probing small-scale structures of the universe. The cold DM leads to an excess of structure on small scales, while warm DM (WDM) or ultra-light DM results in suppressed structure formation on small scales 7, 8. Being sensitive to the small-scale structures, it has been shown that the 21-cm forest can be used to probe the minihalos and dwarf galaxies during the EoR 9, 5, so as to provide precise measurements of small-scale cosmic structures, and can be utilized to measure the mass of DM particles 10.
@@
Our recent work has proposed an effective solution to this problem, namely, by measuring the one-dimensional (1D) power spectrum of the 21-cm forest 11. This approach not only allows for the extraction of faint signals but also breaks the degeneracy between the mass of DM particles and the temperature of the IGM, enabling both to be measured simultaneously. This is because, when transforming from frequency space to $k$-space, the scale-dependence characteristics of the 21-cm forest signal can be displayed, while noise can be effectively suppressed as it lacks such scale dependence, enabling even faint signals to be extracted. Furthermore, the properties of DM particles (whether warm or cold) and the temperature of the IGM mainly affect the shape and amplitude of the power spectrum, respectively. Thus, both can be simultaneously constrained through power spectrum measurements. If this approach can be successfully implemented during the SKA era, it will be significant for research of DM and first galaxies.
@@
To achieve this goal, a series of challenges still need to be overcome. Observationally, it requires a sufficiently sensitive experiment and long observation times to detect the weak signals, as well as a sufficient number of high-redshift radio-bright point sources to suppress the cosmic variance. Additionally, even if we can achieve 21-cm forest observations, it is not straightforward to constrain the properties of DM particles and the cosmic heating history, because currently there is no analytical or empirical model to connect the parameters (such as the mass of DM particles and the temperature of the IGM) with observables (the 1D power spectrum of the 21-cm forest). In this case, in order to achieve parameter inference, we can only rely on a large number of simulations spanning a sufficient large parameter space to calculate observables, and then further construct a likelihood function, thus utilizing Bayesian methods for parameter inference 12. This is how parameter inference is done for constraining DM particles using the Ly-$\alpha$ forest 13. Such a large number of simulations require a significant amount of computing resources, making it both very expensive and extremely time-consuming. Moreover, the computational costs for the 21-cm forest simulations are considerably higher due to its major contributions from smaller scales ($k$ as high as $\sim$100 Mpc-1), compared to the Ly-$\alpha$ forest 14, 2.
@@
Additional challenges arise from the non-Gaussianity of the 21-cm forest signals induced by the nonlinear effect of gravitational clustering 15 and the non-local effect of radiation 16. Due to the non-Gaussian characteristics of the 1D power spectrum, traditional Bayesian methods for parameter inference are somewhat inadequate, potentially leading to severely distorted inference results. Although methods like traditional Gaussian mixture models can fit likelihood functions under non-Gaussian conditions 17, 18, they require significant computational resources, making them less practical for extensive data analysis. Additionally, the correlation between the 1D power spectrum at different scales also affects the precision of parameter inference by coupling errors across scales.
@@
This work proposes a set of solutions to these issues, allowing parameter inference of 21-cm forest observations without the need for an explicit model or the construction of the likelihood function, using only a limited number of simulated samples. This enables us to obtain constraints on the properties of DM particles and the history of cosmic heating efficiently from 1D power spectrum measurements of the 21-cm forest. The key strategy relies on variational inference methods based on deep learning 19, especially the technique of normalizing flows (NFs) 20. These methods have found various applications in fields such as the large-scale generation of neutral hydrogen sky maps 21, 22, parameters inference during the EoR using 21-cm tomography 23, 24, 25, 26, inference of gravitational wave source parameters under the influence of noise transients 27, 28, and the inference of the IGM thermal parameters based on the Ly-$\alpha$ forest 29, 30. Within this framework, not only can a large number of simulated samples be easily generated through generative neural networks, but parameter inference can also be achieved without a likelihood function. This approach addresses the aforementioned challenges of large sample sizes and non-Gaussianity, respectively, for 21-cm forest observations, paving the way for advancing the usage of this probe to constrain the fundamental physics during the cosmic dawn.
@@ Results
We take the same approach as in Ref. 11, and simulate the 21-cm forest signals from the EoR for various X-ray production efficiency of the first galaxies parameterized by $f_{\rm X}$, and for various DM particle masses of $m_{\rm WDM}$ (see Methods). Given the current Hydrogen Epoch of Reionization Array findings, $T_{\rm K}$ is constrained to 15.6 K $<T_{\rm K}<$ 656.7 K at $z\sim 8$ with 95% confidence 31, which demarcates the rough signal level of the 21-cm forest constrained by the heating effect. Therefore, our analysis mainly focus on two characteristic heating levels for discussion: the case with a weaker heating effect ($T_{\rm K}\approx$ 60 K at redshift 9 corresponding to $f_{\rm X}\sim 0.1$) and the case with a stronger heating effect ($T_{\rm K}\approx$ 600 K at redshift 9 corresponding to $f_{\rm X}\sim 1$). For comparison, we also simulate cases with $f_{\rm X}$ = 0 representing an unheated IGM and $f_{\rm X}$ = 3 for an extremely high productivity of X-rays in the early universe. For the DM properties, recent astrophysical observations have put a lower limit on WDM particle mass ($m_{\rm WDM}$) to a few kilo electronvolts (keV) 32, 33, 34, 35, 36, and almost ruled out models below 3 keV 36. Therefore, $m_{\rm WDM}$ ranges from 3 keV to 9 keV in our simulation. Assuming a fiducial WDM model with $m_{\rm WDM}=6$ keV, Fig. 1 shows the signal-to-noise ratio (SNR) of the 1D power spectrum of the simulated 21-cm forest, represented by $P(k)/P^{N}$, with different heating levels for phase-one and phase-two low-frequency array of the SKA (SKA1-LOW and SKA2-LOW), respectively. Our results indicate that a lower heating level ($T_{\rm K}=60~{}\rm K$ at $z=9$) allows for measurement of the 1D power spectrum of the 21-cm forest with an integration time of 100 h per background source using the SKA1-LOW. In contrast, a higher heating level ($T_{\rm K}=600~{}\rm K$ at $z=9$) necessitates 200 h on each source using SKA2-LOW.
Given the high dynamic range required for the 21-cm forest simulations, the simulation-based Bayesian methods for parameter inference are inefficient and almost infeasible, considering the requirement of computational time and resources. To achieve parameter inference with minimal computational expenditure, we first simulate the 21-cm forest signals for a limited sampling of the parameter space. However, the unevenness in simulated data will impair likelihood-free inference methods, potentially leading to inaccurate probability estimations within certain parameter spaces presented in Supplementary Fig. 0.5. To address this, we employ data augmentation using the generative normalizing flow (GNF) to expand the coverage of our datasets beyond the simulated parameter space, thus improving the accuracy of parameter inference results of the subsequent inference normalized flow (INF). The flow chart of our methodology is presented in Fig. 2.
@@
To guarantee a robust parameter inference, we first assess the reliability of the 1D power spectrum of the 21-cm forest generated by the GNF. Figure 3 compares the mean value and standard deviation of the 1D power spectra generated by the GNF with those from simulations. The results confirm that the GNF consistently reproduces the 1D power spectra generated from the training set as presented in the middle panels of Fig. 3. In addition, it effectively generates the 1D power spectra for both interpolated and extrapolated parameter regions as presented in the left and right panels of Fig. 3, demonstrating its generalization capability. To quantify the reasonableness of the GNF model, we applied the coefficient of determination ($R^{2}$) to measure the correlation between the 1D power spectra generated by the GNF and those from simulations. Achieving $R^{2}>0.99$ indicates a high reliability of the GNF in reproducing these 1D power spectra. This examination, combining statistical measures and visual inspections, confirms the reliability of the GNF model in generating the 1D power spectrum of the 21-cm forest.
@@
We used the skewness coefficient ($S$) to describe the asymmetry of the 1D power spectrum at each $k$ in Fig. 4. It shows that at smaller scales and with longer integration times, deviations from a Gaussian distribution increase. This trend is attributed to strong nonlinear effects and a higher proportion of sample variance compared to the observational uncertainty. Additionally, we observed gradual deviations from a normal distribution with increasing $m_{\rm WDM}$ and $T_{\rm K}$. These deviations are attributed to the increased number of halos with higher $m_{\rm WDM}$, which significantly alters the density field, and to temperature variations influencing the 1D power spectrum of the 21-cm forest as $T_{\rm K}$ increases. Additionally, we quantified the similarity between the distributions generated by the GNF and the simulation in Supplementary Fig. 0.5 using the Jensen-Shannon divergence (JSD) 37, 38, a symmetric measure of similarity between two probability distributions. Our analysis revealed that the JSD value is less than 0.05, indicating a high degree of similarity between the distributions generated by the GNF and the simulation. As previously noted with JSD and $S$ analyses, these findings further confirm that under different heating levels, the 1D power spectrum of the 21-cm forest observed with an integration time of 100 h using SKA1-LOW and those measured with an integration time of 200 h using SKA2-LOW exhibit Gaussian and non-Gaussian distributions, respectively. This highlights the capability of the GNF in modeling the 1D power spectrum of the 21-cm forest under various astrophysical conditions.
@@
The correlation coefficient $\rho_{ij}$ between different $k$ bins of the 1D power spectrum of the 21-cm forest is also crucial for accurately estimating the uncertainties in parameter inference. Here, $i$ and $j$ refer to different $k$ bins. Ignoring the correlation in the 1D power spectrum between these different $k$ bins can lead to incorrect error estimates. In supplementary Fig. 0.5, we show the values of $\rho_{ij}$ of the pure signal under different conditions. We find that the correlations at small scales are significant, especially in scenarios with strong heating effects. However, considering the impact of SKA thermal noise in actual observations, as shown in supplementary Fig. 0.5, the dominance of thermal noise at small scales reduces the overall correlation strength. Nonetheless, the correlations remain non-negligible.
First, we need to determine the accuracy of the INF inference parameters. Given the lack of tight constraints on $m_{\rm WDM}$ and $T_{\rm K}$, we explored arbitrary parameter results using the INF in Fig. 5. Our results show that the recovery is mostly near perfect. Next, we determine the precision of the parameters inferred by INF. Similar to findings in Refs. 29, 24, we encounter challenges due to the edge effect, which makes inferences at the boundaries of the parameter space unreliable. However, this can be achieved by expanding the coverage of the simulated parameter space. However, within the central regions of the parameter space, our analysis reveals distinct patterns in the relative error $\varepsilon$ of parameter inference for $m_{\rm WDM}$ and $T_{\rm K}$ under various heating levels, as presented in Supplementary Fig. 0.5. Here, the relative error, defined as the precision, is given by $\varepsilon(\xi)=\sigma(\xi)/\xi$, where $\xi$ represents the true parameter ($m_{\rm WDM}$ or $T_{\rm K}$). We find that $T_{\rm K}$ significantly affects the precision of $m_{\rm WDM}$ estimates. Conversely, $m_{\rm WDM}$ has a relatively small impact on the precision of $T_{\rm K}$ estimates. This is because $T_{\rm K}$ influences the 1D power spectrum values across all scales, while $m_{\rm WDM}$ primarily affects the 1D power spectrum values on small scales, as shown in Supplementary Fig. 0.5. Therefore, as presented in Supplementary Fig. 0.5, $T_{\rm K}$ can achieve high precision by relying solely on large-scale information, almost unaffected by $m_{\rm WDM}$. On the other hand, $m_{\rm WDM}$ mainly relies on small-scale and medium-scale information, which is influenced by $T_{\rm K}$. Therefore, the 1D power spectrum of different scales will break the degeneracy, resulting in a significant improvement in the precision of parameter estimates. This dependence highlights the interplay between these parameters and emphasizes the ability of our network to effectively constrain each parameter.
For low heating level conditions, we utilized SKA1-LOW with an integration time of 100 h to examine the constraint results of posterior parameter inference, assuming $m_{\rm WDM}=6$ keV and $T_{\rm K}=60$ K. We compared the precision of the posterior inference results from the INF with the range predicted by the Fisher matrix as presented in Fig. 6. The INF method successfully recovered the posterior distribution of parameters, with the true values lying within the estimated $1\sigma$ confidence region. This demonstrates precision comparable to the predictions of the Fisher matrix. The final average constraint result of the 1D power spectra of the 21-cm forest for all simulations are $m_{\rm WDM}=6.28_{-1.59}^{+1.61}$ keV and $T_{\rm K}=60.17_{-6.66}^{+6.16}$ K. Detailed simulations of the dataset are described in the Methods section. This confirms the reliability of the INF method in achieving precise parameter inference under low heating conditions.
Under high heating level conditions, using SKA2-LOW with an integration time of 200 h ($m_{\rm WDM}=6$ keV, $T_{\rm K}=600$ K), we observed significantly different results. Although both the Fisher matrix and INF methods presented consistent degeneracy directions, the Fisher matrix approach failed to adequately constrain the parameters due to the non-Gaussian characteristic of the 1D power spectrum of the 21-cm forest. This highlights the effectiveness of likelihood-free inference methods like the INF in such scenarios. The final average constraint results of the 1D power spectra for all simulations are $m_{\rm WDM}=6.47_{-1.88}^{+2.04}$ keV and $T_{\rm K}=627.95_{-91.07}^{+94.18}$ K. These results demonstrate that, while the Fisher matrix approach is inadequate under non-Gaussian conditions, the INF method remains effective.

### v2

@@ Abstract
The hyperfine structure absorption lines of neutral hydrogen in high-redshift radio spectra, known as the 21-cm forest, have been demonstrated through simulations as a powerful probe of small-scale structures governed by dark matter (DM) properties and the thermal history of intergalactic medium (IGM). By measuring the one-dimensional power spectrum of the 21-cm forest, parameter degeneracies can be broken, offering key constraints on the properties of both DM and IGM. However, conventional methods are hindered by computationally expensive simulations and a non-Gaussian likelihood function. To overcome these challenges, we propose a deep learning approach that combines generative normalizing flows for data augmentation and inference normalizing flows for parameter estimation, enabling accurate results from minimally simulated datasets. Using mock data from the Square Kilometre Array, we demonstrate the capability of this deep learning-driven approach to generating posterior distributions, providing a robust tool for probing DM and the cosmic heating history.
@@
Liaoning Key Laboratory of Cosmology and Astrophysics, College of Sciences, Northeastern University, Shenyang 110819, China
@@
International Centre for Theoretical Physics Asia-Pacific, University of Chinese Academy of Sciences, Beijing 100049, China
Taiji Laboratory for Gravitational Wave Universe (Beijing/Hangzhou), University of Chinese Academy of Sciences, Beijing 100049, China
@@
MOE Key Laboratory of Data Analytics and Optimization for Smart Industry, Northeast-ern University, Shenyang 110819, China
Introduction
The hyperfine structure transition of neutral hydrogen atoms from the epoch of reionization (EoR) creates narrow absorption lines on spectra of high-redshift radio bright sources (e.g. radio-loud quasars), with a rest-frame wavelength of 21 cm. The 21-cm absorption lines from various structures at different distances along the line of sight are redshifted to different frequencies, giving rise to forest-like features on the spectra, which is called the 21-cm forest 1, 2, 3, 4, 5, 6 in analogy to the Lyman-$\alpha$ (Ly-$\alpha$) forest at lower redshifts. Simulations suggest that the 21-cm forest could serve as a sensitive probe of the cosmic heating history 4, 5. After the formation of the first galaxies, the generated light not only ionizes hydrogen atoms but also heats the gas simultaneously (with X-rays being the primary heating source). As the spin temperature of the gas increases, the absorption at 21 cm weakens (reducing the optical depth), thus causing the 21-cm forest signal to become weaker. In principle, the strength of the signal can be used to measure the cosmic heating history, thereby gaining insight into the properties of the first galaxies.
On the other hand, understanding the properties of dark matter (DM) hinges crucially on probing small-scale structures of the universe. The cold DM (CDM) leads to an excess of structure on small scales, while warm DM (WDM) or ultra-light DM results in suppressed structure formation on small scales 7, 8. Being sensitive to the small-scale structures, it has been shown that the 21-cm forest can be used to probe the minihalos and dwarf galaxies during the EoR 9, 5, so as to provide precise measurements of small-scale cosmic structures, and can be utilized to measure the mass of DM particles 10.
@@ Introduction
Our recent work has proposed an effective solution to this problem, namely, measuring the one-dimensional (1D) power spectrum of the 21-cm forest 11. This approach not only allows for the extraction of faint signals but also breaks the degeneracy between the mass of DM particles and the temperature of the IGM, enabling both to be measured simultaneously. This is because, when transforming from frequency space to $k$-space, the scale-dependence characteristics of the 21-cm forest signal can be displayed, while noise can be effectively suppressed as it lacks such scale dependence, enabling even faint signals to be extracted. Furthermore, the properties of DM particles (whether warm or cold) and the temperature of the IGM mainly affect the shape and amplitude of the power spectrum, respectively. Thus, both can be simultaneously constrained through power spectrum measurements. If this approach can be successfully implemented during the SKA era, it will be significant for research of DM and first galaxies.
@@
To achieve this goal, a series of challenges still need to be overcome. Observationally, it requires a sufficiently sensitive experiment and long observation times to detect the weak signals, as well as a sufficient number of high-redshift radio-bright point sources to suppress the cosmic variance 12. Additionally, even if we can achieve 21-cm forest observations, it is not straightforward to constrain the properties of DM particles and the cosmic heating history, because currently there is no analytical or empirical model to connect the parameters (such as the mass of DM particles and the temperature of the IGM) with observables (the 1D power spectrum of the 21-cm forest). In this case, to achieve parameter inference, we can only rely on a large number of simulations spanning a sufficient large parameter space to calculate observables, and then construct a likelihood function, thus utilizing Bayesian methods for parameter inference 13. This is how parameter inference is done for constraining DM particles using the Ly-$\alpha$ forest 14. Such a large number of simulations require a significant amount of computing resources, making it both very expensive and extremely time-consuming. Moreover, the computational costs for the 21-cm forest simulations are considerably higher due to its major contributions from smaller scales ($k$ as high as $\sim 100$ Mpc-1), compared to the Ly-$\alpha$ forest 15, 2.
@@
Additional challenges arise from the non-Gaussian likelihood function of the 21-cm forest signals induced by the nonlinear effect of gravitational clustering 16 and the non-local effect of radiation 17. Due to the non-Gaussian likelihood function for the 1D power spectrum, traditional Bayesian methods for parameter inference are somewhat inadequate, potentially leading to severely distorted inference results. Although methods like traditional Gaussian mixture models can fit likelihood functions under non-Gaussian conditions 18, 19, they require significant computational resources, making them less practical for extensive data analysis. Additionally, the correlation between the 1D power spectrum at different scales also affects the precision of parameter inference framework, as the coupling of errors across scales increases the complexity of modeling the likelihood.
@@
This work proposes a set of solutions to these issues, allowing parameter inference of 21-cm forest observations without the need for an explicit model or the construction of the likelihood function, using only a limited number of simulated samples. This enables us to obtain constraints on the properties of DM particles and the history of cosmic heating efficiently from 1D power spectrum measurements of the 21-cm forest. The key strategy relies on variational inference methods based on deep learning 20, especially the technique of normalizing flows (NFs) 21. These methods have found various applications in fields such as the large-scale generation of neutral hydrogen sky maps 22, 23, parameter inference during the EoR using 21-cm tomography 24, 25, 26, 27, inference of gravitational wave source parameters under the influence of noise transients 28, 29, and the inference of the IGM thermal parameters based on the Ly-$\alpha$ forest 30, 31. Within this framework, not only can a large number of simulated samples be easily generated through generative neural networks, but parameter inference can also be achieved without a likelihood function. This approach addresses the aforementioned challenges of large sample sizes and a non-Gaussian likelihood function, respectively, for 21-cm forest observations, paving the way for advancing the usage of this probe to constrain the fundamental physics during the cosmic dawn.
@@ Results
We take the same approach as Ref. 11, and simulate the 21-cm forest signals from the EoR for various X-ray production efficiency of the first galaxies parameterized by $f_{\rm X}$, and for various DM particle masses of $m_{\rm WDM}$ (see Methods). The parameter $f_{\rm X}$ controls the efficiency of X-ray heating in the early universe, directly influencing the IGM temperature $T_{\rm K}$. Given the current Hydrogen Epoch of Reionization Array (HERA) findings, $T_{\rm K}$ is constrained to 15.6 K $<T_{\rm K}<$ 656.7 K at $z\sim 8$ with 95% confidence 32, which demarcates the rough signal level of the 21-cm forest constrained by the heating effect. Therefore, our analysis mainly focuses on two characteristic heating levels for discussion: the case with a weaker heating effect ($T_{\rm K}\approx$ 60 K at redshift 9 corresponding to $f_{\rm X}\sim 0.1$) and the case with a stronger heating effect ($T_{\rm K}\approx$ $600$ K at redshift 9 corresponding to $f_{\rm X}\sim 1$). As $f_{\rm X}$ increases, $T_{\rm K}$ rises, regulating the X-ray heating in the early universe and strongly influencing the detectability of the 21-cm forest signal. For comparison, we also simulate cases with $f_{\rm X}=0$ representing an unheated IGM and $f_{\rm X}=3$ for extremely high productivity of X-rays in the early universe. For the DM properties, although some recent astrophysical observations suggest a lower limit on the WDM particle mass ($m_{\rm WDM}$) of approximately 6 kilo electronvolts (keV), models with lower WDM particle mass have not been definitively excluded 33, 34, 35, 36, 37. Furthermore, as the value of $m_{\rm WDM}$ increases, the structure formation characteristics of WDM models on small scales become increasingly similar to those of CDM model. We therefore simulate WDM models with mass in the several keV range for subsequent analysis. Assuming a fiducial WDM model with $m_{\rm WDM}=6$ keV, Figure 1 shows the signal-to-noise ratio (SNR) of the 1D power spectrum of the simulated 21-cm forest, represented by $P(k)/P^{\rm N}$, where $P(k)$ is the 1D power spectrum of the pure 21-cm forest signal and $P^{\rm N}$ is thermal noise of the 1D power spectrum, with different heating levels for the noise levels corresponding to phase-one and phase-two low-frequency arrays of the SKA (SKA1-LOW 38 and SKA2-LOW 39, 40), respectively. Our results indicate that a lower heating level ($T_{\rm K}=60~{}\rm K$ at $z=9$) allows for measurement of the 1D power spectrum of the 21-cm forest with an integration time of $100$ h per background source with a flux density of $10$ mJy using the SKA1-LOW. In contrast, a higher heating level ($T_{\rm K}=600~{}\rm K$ at $z=9$) necessitates 200 h on each source with the same flux density using SKA2-LOW.
Given the high dynamic range required for the 21-cm forest simulations, the simulation-based Bayesian methods for parameter inference are inefficient and almost infeasible, considering the requirement of computational time and resources. For instance, simulating the density field for $m_{\rm WDM}=6$ keV requires approximately $600$ h on $64$ cores (Intel(R) Xeon(R) Gold 6271C CPU @ 2.60GHz). To achieve parameter inference with minimal computational expenditure, we first simulate the 21-cm forest signals for a limited sampling of the parameter space. Specifically, $m_{\rm WDM}$ ranges from $3$ keV to $9$ keV with intervals of $1$ keV, while $T_{\rm K}$ ranges from $40$ K to $80$ K with intervals of $5$ K for weaker heating levels, and from $400$ K to $800$ K with intervals of $100$ K for stronger heating levels. This parameter space results in a dataset of $567000$ samples (corresponding to $7$ values of $m_{\rm WDM}$ and $9$ values of $T_{\rm K}$, for a total of $7\times 9\times 9000=567000$ samples; see Methods) for weaker heating levels, which, while sufficiently large, is unevenly distributed and does not fully cover the entire parameter space. To ensure robustness, the validation dataset, split from the training data, is employed to fine-tune model parameters and prevent overfitting. The testing dataset, also separated from the initial dataset, is reserved exclusively for evaluating the model’s predictive performance on unseen data.
As shown in Supplementary Fig. 0.8, this non-uniform sampling of the parameter space impairs likelihood-free inference methods, leading to inaccurate probability estimations in under-sampled regions of the parameter space. However, as presented in Supplementary Fig. 0.8, while the parameter space distribution is irregular, the corresponding power spectrum space exhibits continuity, indicating that the power spectra are less sensitive to gaps in parameter sampling. To leverage this property, our framework combines a generative normalizing flow (GNF) to fill the gaps in the parameter space and an inference normalizing flow (INF) to accurately infer parameters, as shown in Fig. 2.
0.2 Data augmentation.
@@
To guarantee a robust parameter inference, we first assess the reliability of the 1D power spectrum of the 21-cm forest generated by the GNF. Figure 3 compares the mean value and standard deviation of the 1D power spectra generated by the GNF with those from simulations. The results confirm that the GNF consistently reproduces the 1D power spectra generated from the training set as presented in (b, e, h) of Fig. 3. In addition, it effectively generates the 1D power spectra for both interpolated and extrapolated parameter regions as presented in (a, d, g) and (c, f, i) of Fig. 3, demonstrating its generalization capability. To quantify the reasonableness of the GNF model, we applied the coefficient of determination ($R^{2}$) to measure the correlation between the 1D power spectra generated by the GNF and those from simulations. Achieving $R^{2}>0.99$ indicates a high reliability of the GNF in reproducing these 1D power spectra. This examination, combining statistical measures and visual inspections, confirms the reliability of the GNF model in generating the 1D power spectrum of the 21-cm forest. Additionally, using the GNF, approximately $24000$ 1D power spectra can be generated in about $1$ s, a speed unattainable by traditional simulations. In contrast, generating the same number of samples with traditional simulations would require approximately $280$ h.
@@
We used the skewness coefficient ($S$) to describe the asymmetry of the 1D power spectrum at each $k$ in Fig. 4. It shows that at smaller scales and with longer integration times, deviations from a Gaussian distribution increase. This trend is attributed to the increasing dominance of small-scale strong nonlinear effects compared to thermal noise. Additionally, we observed gradual deviations from a normal distribution with increasing $m_{\rm WDM}$ and $T_{\rm K}$. These deviations are attributed to the increased number of halos with higher $m_{\rm WDM}$, which significantly alters the density field, and to temperature variations influencing the 1D power spectrum of the 21-cm forest as $T_{\rm K}$ increases. Additionally, we quantified the similarity between the distributions generated by the GNF and the simulation in Supplementary Fig. 0.8 using the Jensen-Shannon divergence (JSD) 41, 42, a symmetric measure of similarity between two probability distributions. Our analysis revealed that the JSD value is less than $0.05$, indicating a high degree of similarity between the distributions generated by the GNF and the simulation. As previously noted with JSD and $S$ analyses, these results confirm that under different heating levels, the 1D power spectrum of the 21-cm forest observed with an integration time of $100$ h using SKA1-LOW and those measured with an integration time of $200$ h using SKA2-LOW exhibit Gaussian and non-Gaussian distributions, respectively. This highlights the capability of the GNF in modeling the 1D power spectrum of the 21-cm forest under various astrophysical conditions.
@@
The correlation coefficient $\rho_{ij}$ between different $k$ bins of the 1D power spectrum of the 21-cm forest is also crucial for accurately estimating the uncertainties in parameter inference. Here, $i$ and $j$ refer to different $k$ bins. Ignoring the correlation in the 1D power spectrum between these different $k$ bins can lead to incorrect error estimates. In Supplementary Fig. 0.8, we show the values of $\rho_{ij}$ of the pure signal under different conditions. We find that the correlations at small scales are significant, especially in scenarios with strong heating effects. However, considering the impact of SKA thermal noise in actual observations, as shown in Supplementary Fig. 0.8, the dominance of thermal noise at small scales reduces the overall correlation strength. Nonetheless, the correlations remain non-negligible. Therefore, although combining GNF with Bayesian inference allows us to obtain posterior distributions, the aforementioned correlations significantly increase the complexity of modeling the likelihood function. As a result, generating $2^{16}$ posterior samples requires several days of computation on $32$ cores.
0.3 Parameter inference.